If three coins are tossed, what are the mean, variance, and standard deviation of the number of heads that occur?


1
Expert's answer
2022-01-20T13:45:59-0500

We assume that the probabilities of getting heads and tails are the same

"p = q = \frac{1}{2}"

Using the Bernoulli formula, we find the probabilities that there will be 0, 1, 2, 3 and 4 heads, respectively

"P\left( {x = 0} \right) = {q^3} = \frac{1}{{8}}"

"P(x = 1) = C_3^1p{q^2} = \frac{3}{{8}}"

"P(x= 2) = C_3^2{p^2}{q} = \frac{3}{{8}}"

"P\left( {x = 3} \right) = {p^3} = \frac{1}{{8}}"

We have a distribution series


So, the mean is

"M(x) = {\sum x _i}{p_i} = \frac{{0 \cdot 1 + 1 \cdot 3 + 2 \cdot 3 + 3 \cdot 1}}{8} = \frac{{12}}{8} = \frac{3}{2}"

The variance is

"V(x) = M\left( {{x^2}} \right) - {M^2}(x) = \frac{{0 \cdot 1 + 1 \cdot 3 + 4 \cdot 3 + 9 \cdot 1}}{8} - \frac{9}{4} = \frac{3}{4}"

standard deviation is

"\sigma \left( x \right) = \sqrt {V(x)} = \frac{{\sqrt 3 }}{2}"

Answer: "M(x) = \frac{3}{2}" ; "V(x) = \frac{3}{4}" ; "\sigma \left( x \right) = \frac{{\sqrt 3 }}{2}"